OpenStudy (anonymous):

Charlie has a piece of toast that has cream cheese on one side, and he dropped it once. One time, it landed with the cream cheese side up. If he drops it two more times, what is the probability that it will have landed cream cheese side up a total of three times? one over four one over three one over two three over four

OpenStudy (mathstudent55):

Is the landing of the toast cream cheese side up or down a completely random thing? In other words, is there as much chance of the toast landing cream cheese side up or cream cheese side down?

OpenStudy (anonymous):

its completely random and no

OpenStudy (anonymous):

first you have to take in the gravitational rotation that takes place as the toast is falling since there's only cream cheese on one side , then you have to solve all the mathematical equations to figure out how much cream cheese fell off the first time the toast fell . And after you've done all the math you discover that the correct answer is B , one over three

zepdrix (zepdrix):

I think Charlie needs to stop eating food off of the floor. But in all seriousness, see how the `first event` has already happened? We don't use that to calculate our odds. So we have two events that need to happen. What are the odds of landing cream cheese side up one more time? One in two, yes? 50% chance. This needs to happen `two more times`. So we multiply these values together.\[\Large\rm \frac{1}{2}\cdot\frac{1}{2}\]Understand what I was saying?

OpenStudy (anonymous):

yes sir sorry and hahahahah I agree... food candy is not as good as it sounds :p

OpenStudy (mathstudent55):

If there is as much chance of the toast landing cream cheese side up or down, then these are all independent events with a 50-50 probability of the result in each toss. Possible results for the next two tosses are (U = cream cheese up; D = cream cheese down): U U U D D U D D The outcome of interest is U U, which is 1/4 of all outcomes for the last two tosses. Since the first outcome of U is already done, it does not influence the final result.